Question: Find the common solution of y=x^2-1 and x+y=1, graphically. Thank you for your help!
Answer by mananth(16949)

y=x^2-1 and x+y=1
y=x^2-1
this is a parabola
generate values of x& y
x=0, y=-1 (0,-1)
y=0,x=1 (1,0)
x=-2, y=3 (-2,3)
x=2,y=3 (2,3)
and x+y=1
x=0,y=1 (0,1)
y=0,x=1(1,0)
Plot the points draw the curve and the straight line

The solutions are (1,0) , (-2,3)
m.ananth@hotmail.ca